OpenStudy (anonymous):

Simplify the expression 12 divide 4 + 2

OpenStudy (solomonzelman):

12 ÷ 4 + 2 3 + 2 your answer is ?

OpenStudy (anonymous):

10?

OpenStudy (cj49):

4+2=6 12/6=2

OpenStudy (anonymous):

So 8?

OpenStudy (cj49):

Simplify the expression 12 divide 4 + 2

OpenStudy (cj49):

this is the question ryt

OpenStudy (anonymous):

Yes

OpenStudy (cj49):

then it wud be 12/6=2 the ans is 2